OpenStudy (anonymous):

How many possible phone numbers are there in a city that has two area codes? A.) 180,000 B.) 9,565,938 C.) 18,000,000 D.) 20,000,000

OpenStudy (31356):

The only differences in the phone numbers are the last 3 digits. How many possible phone numbers can be assigned using this system? 10 * 10 * 10 = 1000 ============================== Many large cities have run out of phone numbers and so a new area code must be introduced. How many different phone numbers are there in a single area code if the first digit cannot be zero? (area code) _ _ _ - _ _ _ _ First digit cannot be zero (must be 1-9) All other digits can be 0-9 9 * 10 * 10 * 10 * 10 * 10 * 10 = 9,000,000 ============================== How many different phone numbers are possible using a 3-digit area code and a 7-digit phone number if the first digit of the area code and phone number cannot be zero? _ _ _ - _ _ _ - _ _ _ _ ^ ........ ^ : Digits 1 through 9 All others: Digits 0 through 9 9 * 10 * 10 * 9 * 10 * 10 * 10 * 10 * 10 * 10 = 8,100,000,000 ============================== A shipping service offers to send packages by ground delivery using 2 different companies, by next day air using 3 different companies, and by 2-day air using 3 different companies. How many different shipping options does the company offer? Since all methods of delivery are independent, just add them (i.e. you will NOT be using ground delivery AND 2-day air delivery, or next day air delivery AND 2-day air delivery). 2 + 3 + 3 = 8

OpenStudy (31356):

Medal?

OpenStudy (31356):

Each area code has about 7,800,000

OpenStudy (the_fizicx99):

Seriously? http://answers.yahoo.com/question/index?qid=20110506135531AAHVMGG

OpenStudy (31356):

That's my works cited page.

OpenStudy (31356):

I was going to post it; so others wouldn't be calling it (plagiarism)!

OpenStudy (the_fizicx99):

You're implying that you wrote that on yahoo?